OpenStudy (anonymous):

Given: ABC is similar to XYC. Enter answer as a simplified, improper fraction. If BY = 4, YC = 7, XC = 10, then AC

OpenStudy (anonymous):

I think is is 9/2

OpenStudy (anonymous):

\[AC = \frac{110}{7}\]BC=7+4=11 AC=AX+10 \[\frac{10}{7}=\frac{AX+10}{11}\]cross and time 10*11=7AX+70 110 =7AX+70 AX =(110-70)/7 AX =40/7 AC= AX +10 \[AC= \frac{40}{7}+10\] deniminator is 7 \[ AC =\frac{40+70}{7} \] \[AC= \frac{110}{7}\] if u want decimal u caculater

OpenStudy (anonymous):

If I deserve medal then give me one